OpenStudy (marsmcnugget):

3. Joe solved and justified each step as shown. Identify Joe’s error and fix his mistake. Step Justification Given equation Multiplication Property of Equality Subtraction Property of Equality

OpenStudy (marsmcnugget):

joe solved 3+ x/2 =10

OpenStudy (marsmcnugget):

step justification 3+ x/2 =10 given equation x/2 =7 multiplication property of equality x=14 subtraction property of equality